# On Efficient Variants of Segment Anything Model: A Survey

### v3

@@ 3.1.2 Knowledge Distillation based Methods
Figure 8: The framework of the encoder-only distillation for MobileSAM [43]. The small ViT-based image encoder, TinyViT [114], is distilled from SAM’s ViT-H and the prompt-guided mask decoder is inherited from SAM directly.
@@
Zhang et al. [43] made the first attempt to replace SAM’s heavy ViT encoder with the more efficient TinyViT [114], resulting in the integrated model named MobileSAM . As highlighted in [19], training SAM from scratch requires multiple days and 128 GPUs. MobileSAM attributes this complexity to the challenge of optimizing both the encoder and decoder simultaneously. To address this, they propose an encoder-only distillation strategy as illustrated in Fig. 8, which aims to transfer the visual representation capabilities of ViT-H to TinyViT. The loss function used is a simple Mean Square Error (MSE) between the output image embeddings of the two encoders. Further fine-tuning of the prompt encoder or mask decoder is optional and can lead to improved accuracy. The image encoder of MobileSAM has been further applied to HQ-SAM [104], a variant significantly boosting SAM’s mask quality, for mobile efficiency. The resulted model, namely light HQ-SAM, exceeds MobileSAM in performance with negligible gap in inference speed, proving it an excellent SAM variant trading off well between accuracy and efficiency.
@@
Similar to MobileSAM, the later proposed ESAM [121] utilizes EfficientFormerV2 [120] as its backbone, aiming for improved performance in CPU environments, particularly resource-constrained medical devices. Given that expert models [177, 178, 179] often outperform SAM in medical applications, ESAM incorporates a novel Knowledge Distillation (KD) strategy called Holistic Knowledge Distillation (HKD) to transfer knowledge from an expert model to ESAM. HKD involves two components: distillation on feature maps and distillation on output masks. For the feature map distillation, three different methods with varying focuses [180, 181, 182] are combined to guide the learning process. For the output mask distillation, ESAM uses the Mean Square Error (MSE) loss between the teacher’s and student’s masks, supplemented by a Binary Cross-Entropy (BCE) loss between the teacher’s masks and the ground truth masks. A Teacher Guided Module (TGM) is introduced, to further align the feature maps between the expert model and ESAM.
@@
Figure 9: The framework of the hard mining full-stage knowledge distillation for TinySAM [44]. It contains two novel strategies to improve the quality of distillation: the Hard Mask Weighting and the Hard Prompt Sampling.
@@
Not long after MobileSAM’s publishing, Shu et al. [44] conducted an analysis on it, identifying that encoder-only distillation can lead to significant performance degradation. To address this issue, they propose the Hard Mining Full-Stage Knowledge Distillation strategy for more effective distillation, as illustrated in Fig. 9. With a structure identical to MobileSAM, they developed a new SAM variant called TinySAM by training with this improved KD strategy. Specifically, the strategy supervises not only the image embeddings but also the output tokens and output masks, all using L1 Loss. To further enhance the distillation process, they introduced the Hard Mask Weighting strategy, which assigns larger weights to masks that are harder to predict, thereby improving learning efficiency. The factor $H$ is calculated as follows,
@@
Researchers from NVIDIA introduced a new SAM variant, NanoSAM [138], based on MobileSAM. It is designed to achieve real-time performance on NVIDIA Jetson Orin platforms using NVIDIA TensorRT. NanoSAM replaces the ViT-based encoder with a pure convolutional network, specifically ResNet18 [112], while retaining the other components from MobileSAM. NanoSAM is distilled from MobileSAM, and both models are retrained with TensorRT for optimized performance. The image encoder of MobileSAM is optimized with FP32 precision, whereas NanoSAM’s image encoder is built using FP16 precision for faster execution. Inference latency results on Jetson Orin Nano and Jetson AGX Orin demonstrate that NanoSAM is 5x faster than MobileSAM, with minimal accuracy loss. Very recently, a similar work, which builds up an ultra light-weight variant based on the U-Net architecture, named PicoSAM2 [151], has successfully deployed SAM on Sony IMX500 (a camera sensor with 8MB memory limit.
@@
Due to MobileSAM with TinyViT as backbone still encounters challenges on resource-constrained devices, Wang et al. [113] developed an novel efficient SAM variant, RepViT-SAM, using their newly proposed CNN-based backbone, RepViT [45], as the image encoder. The core idea behind RepViT is to integrate the effective design principles of efficient Vision Transformers (ViTs) into lightweight CNNs. These design principles are applied at three levels: block-level, macro-level, and micro-level. At the block level, RepViT separates the token mixer and channel mixer [183], reduces the expansion ratio [184], and increases the width of blocks. For the macro design, it incorporates early convolutions [185] as the input stem, deepens the down-sampling layers [186], employs a simpler classifier, and adjusts the block ratios across stages [187]. At the micro level, only 3x3 convolutions are used, and Squeeze-and-Excitation layers [188] are applied only in odd-numbered blocks. RepViT-SAM is trained by following the same distillation pipeline in [43], claiming to achieve a 10x increase in inference speed compared to MobileSAM.
@@
Figure 10: The framework of distilling EdgeSAM [46]. It consists of the encoder-only distillation and the prompt-in-the-loop distillation.
@@
Concurrent with the development of RepViT-SAM, Zhou et al. [46] observe that MobileSAM [43] still struggles to achieve real-time performance when deployed on edge devices, such as mobile phones. To address this, they introduced EdgeSAM, which replaces the Transformer-based encoder with the more lightweight and efficient, pure CNN-based RepViT, aiming for better performance on resource-constrained devices. Similar to the approach in [121], Zhou et al. argue that encoder-only distillation is inadequate, as it is task-agnostic and does not fully capture the model’s task-specific needs. To overcome this, they propose the Prompt-in-the-Loop Distillation method, which adds additional supervision to the output masks. The ”Prompt-in-the-Loop” refers to a dynamic sampling strategy that iteratively samples new prompts from the non-overlapping zones of the teacher’s and student’s predicted masks. After several iterations, the accumulated loss is backpropagated to update both the encoder and decoder. To further enhance the output quality, EdgeSAM offers an optional module that embeds granularity priors from specific datasets. The overall framework for distilling EdgeSAM is illustrated in Fig. 10.
@@
To further improve the efficiency of backbone while maintaining strong capability of feature extraction, Zhang et al. [122] introduced EfficientViT [116] as image encoder, building up the powerful variant, EfficientViT-SAM. The primary advantage of EfficientViT is its use of the ReLU linear attention mechanism, which facilitates global information interaction while enhancing hardware efficiency. By eliminating the hardware-unfriendly softmax operations and replacing them with ReLU, the attention calculation is reformulated as follows,
@@
Following the similar idea of applying high-efficiency attention mechanisms, Shen et al. [128] proposed FastSAM3D, an efficient segment-anything model specifically designed for 3D volumetric medical images. The key contribution of this work is the development of the 3D Sparse Flash Attention mechanism. This novel attention approach combines the advantages of the 3D Dilated Attention [189], which extends the receptive field, with FlashAttention [127] to accelerate computations. FastSAM3D utilizes a modified ViT-Tiny as the image encoder, distilled from the ViT-Base encoder, ensuring efficiency without compromising performance. The authors implemented a layer-wise progressive distillation strategy to iteratively align feature maps between the two encoders, as shown in Fig. 11.
Figure 11: The framework of layer-wise progressive distillation of FastSAM3D [128]. It takes the ViT-B as teacher model to distill the efficient image encoder, a 6-layer ViT-Tiny variant.
Building on the approach of FastSAM3D, Song et al. [129] introduced a 2D variant called SAM-Lightening. Like its 3D counterpart, SAM-Lightening combines Sparse/Dilated Attention and FlashAttention as a replacement for the standard attention mechanism, while retaining the same image encoder. The key difference lies in the Knowledge Distillation (KD) strategy, referred to as Dynamic Layer-Wise Distillation (DLD). In DLD, a series of time-varying weights $\in[0,1]$ are applied to determine which layers need to be updated and how much each layer contributes to the update process. As training progresses, the entire architecture is gradually optimized. SAM-Lightening reportedly achieves a 30× acceleration compared to SAM-H. The overall distillation framework is similar to FastSAM3D, with the main change being the substitution of the distillation strategy for dynamic layer-wise distillation.
A recent work by Yuan et al. [136], RWKV-SAM, represents a significant advancement in accelerating SAM by introducing the popular linear attention model [133, 117] as an efficient backbone. In their study, they compare RWKV-based and Mamba-based architectures and select the RWKV-based approach to construct a lightweight version of SAM. The backbone is a hybrid design, with the first two stages consisting of Mobile Convolution Blocks from [119], and the final stage built using Vision RWKV blocks [135]. More details on RWKV can be found in Section 2.2.2. Additionally, a refinement module is incorporated into the SAM-like architecture to enhance mask quality by fusing features from different levels generated at each stage. The model is trained using a ”distillation-finetune” strategy, where knowledge from SAM-H is first distilled into the backbone, followed by fine-tuning of the entire model. RWKV-SAM demonstrates significant improvements in efficiency, while maintaining comparable segmentation performance to SAM.
@@ 3.1.3 Quantization based Methods
As detailed in previous section, TinySAM [44] leverages the Hard Mining Full-Stage Distillation to improve the effectiveness of transferring knowledge. To further contract the scale of TinySAM, Shu et al. adopt a post-training quantization to the encoder of TinySAM and the quantified version is named as Q-TinySAM. They take the quantization basically following the instructions in [145]. For the matrix multiplication in ViT, $O=AB$ would be quantified to $\hat{O}=\hat{A}\hat{B}$, with scaling factors $s_{A},s_{B}$. The researchers takes a alternative and iterative search strategy for the best $s_{A},s_{B}$ which can minimize the distance between $O$ and $\hat{O}$. The hessian guided metric is used to measure the distance which is approximately calculated as follows:
@@
Instead of taking quantization in an encoder-only manner used for Q-TinySAM, Lv et al. [47] propose a novel framework that can directly take post-training quantization on SAM, namely PTQ4SAM. They start the work by first discovering the two challenges after traditional PTQ: 1) the existence of bimodal distribution which has negative effects on quantization quality and 2) the obvious discrepancy in the distribution of different attention mechanisms. Therefore, researchers bring up two strategies to solve them respectively: the Bimodal Integration (BIG) and the Adaptive Granularity Quantization (AGQ). For the Bimodal Integration, a sign factor $\gamma$ is introduced to convert the bimodal distribution into a normal distribution. For the Adaptive Granularity Quantization, the key is to adopt an adaptive parameter $\tau$ to adjust the base in the Log2 quantizer [190]. With the use of the AGQ strategy, the discrepancy between post-softmax distributions of different attention is expected to narrow down effectively. The PTQ4SAM is a plug-and-play SAM variant that can be easily deployed to downstream tasks.
Concurrent with PTQ4SAM, a counterpart named PQ-SAM [152], further exploits the post-training quantization method customized for SAM’s activations, achieving the first usable 4-bit PTQ for SAM. The core idea lies in transferring SAM’s highly asymmetric activation distribution with extreme outliers into quantization-friendly distribution, for which the Grouped Activation Distribution Transformation (GAPT) is proposed. GAPT firstly adopts tensor-wise outlier truncating to eliminate outliers in activation values, followed by the channel-wise outlier grouping to classify channels of truncated tensors into clusters based on their range. Pairs of learnable shifting and scaling sizes are subsequently allocated to each channel group, reshaping the distribution as follows:
@@ 3.1.4 Pruning based Methods
Figure 12: The overall pipeline of the alternate slimming strategy for SlimSAM [48]. It takes two steps to contract the heavy image encoder of SAM into a light-weight one: 1) Embedding pruning and bottleneck aligning; 2) Bottleneck pruning and embedding aligning.
Chen et al. [48] were the first to develop an effective pruning strategy for reducing the size and complexity of SAM, resulting in the model known as SlimSAM. As discussed in Section 2.3.3, pruning algorithms aim to remove redundant parameters either structurally or individually. When applied to SAM’s heavy encoder, the initial step involves estimating the importance of weights and activation values to determine which should be pruned. The core idea behind assessing importance is to evaluate the discrepancy in loss generated with and without the given parameter. SlimSAM introduces the Disturbed Taylor Importance method, which uses the first-order Taylor expansion to approximate the importance of parameters and introduces Gaussian Noise $\mathcal{N}$ to prevent gradients from becoming zero. This process is formulated as follows,
@@
Unlike Chen et al. [48] adopting the pruning and post-alignment strategy to lighten SAM in a straightforward manner, Abebe et al. [153] incorporates pruning methods with the one-shot Neural Architecture Search [191] (NAS, an effective technique of automating the design of efficient neural networks), building up a weight-sharing supernetwork of SAM, named SuperSAM. The overall architecture of SuperSAM is illustrated as Fig. 13. SuperSAM works as a search space of SAM-like subnetworks with size 30-70% smaller than original SAM-B, while some of them still maintaining considerable performance are considered as efficient variants of SAM. The supernetwork is trained by sampling and training subnetworks with flexible image encoder, in which the structured pruning is applied to remove insignificant Transformer layers, with further weight reordering and slicing of MLP in remaining layers. Once the SuperSAM is built up, techniques such as OpenTuner [192] can be deployed to search for the architecture meeting both efficiency and accuracy requirement. The emerge of SuperSAM might further broaden the domain of investigating efficient SAM variants.
Figure 13: The architecture of SuperSAM [153] consists of a flexible image encoder and SAM-like prompt-guided mask decoder. Lightweight SAM variants can be built up by sampling appropriate image encoder meeting predefined efficiency and accuracy requirements.

### v4

@@ 3.1.2 Knowledge Distillation based Methods
Figure 8: The framework of the encoder-only distillation for MobileSAM [zhang2023faster]. The small ViT-based image encoder, TinyViT [wu2022tinyvit], is distilled from SAM’s ViT-H and the prompt-guided mask decoder is inherited from SAM directly.
@@
Zhang et al. [zhang2023faster] made the first attempt to replace SAM’s heavy ViT encoder with the more efficient TinyViT [wu2022tinyvit], resulting in the integrated model named MobileSAM . As highlighted in [kirillov2023segment], training SAM from scratch requires multiple days and 128 GPUs. MobileSAM attributes this complexity to the challenge of optimizing both the encoder and decoder simultaneously. To address this, they propose an encoder-only distillation strategy as illustrated in Fig. 8, which aims to transfer the visual representation capabilities of ViT-H to TinyViT. The loss function used is a simple Mean Square Error (MSE) between the output image embeddings of the two encoders. Further fine-tuning of the prompt encoder or mask decoder is optional and can lead to improved accuracy. The image encoder of MobileSAM has been further applied to HQ-SAM [ke2024segment], a variant significantly boosting SAM’s mask quality, for mobile efficiency. The resulted model, namely light HQ-SAM, exceeds MobileSAM in performance with negligible gap in inference speed, proving it an excellent SAM variant trading off well between accuracy and efficiency.
@@
Similar to MobileSAM, the later proposed ESAM [Zhao2023ESAM] utilizes EfficientFormerV2 [li2023rethinking] as its backbone, aiming for improved performance in CPU environments, particularly resource-constrained medical devices. Given that expert models [wu2023medical, wang2022stepwise, chang2023esfpnet] often outperform SAM in medical applications, ESAM incorporates a novel Knowledge Distillation (KD) strategy called Holistic Knowledge Distillation (HKD) to transfer knowledge from an expert model to ESAM. HKD involves two components: distillation on feature maps and distillation on output masks. For the feature map distillation, three different methods with varying focuses [chen2021distilling, shu2021channel, yang2022focal] are combined to guide the learning process. For the output mask distillation, ESAM uses the Mean Square Error (MSE) loss between the teacher’s and student’s masks, supplemented by a Binary Cross-Entropy (BCE) loss between the teacher’s masks and the ground truth masks. A Teacher Guided Module (TGM) is introduced, to further align the feature maps between the expert model and ESAM.
@@
Figure 9: The framework of the hard mining full-stage knowledge distillation for TinySAM [shu2023tinysam]. It contains two novel strategies to improve the quality of distillation: the Hard Mask Weighting and the Hard Prompt Sampling.
@@
Not long after MobileSAM’s publishing, Shu et al. [shu2023tinysam] conducted an analysis on it, identifying that encoder-only distillation can lead to significant performance degradation. To address this issue, they propose the Hard Mining Full-Stage Knowledge Distillation strategy for more effective distillation, as illustrated in Fig. 9. With a structure identical to MobileSAM, they developed a new SAM variant called TinySAM by training with this improved KD strategy. Specifically, the strategy supervises not only the image embeddings but also the output tokens and output masks, all using L1 Loss. To further enhance the distillation process, they introduced the Hard Mask Weighting strategy, which assigns larger weights to masks that are harder to predict, thereby improving learning efficiency. The factor $H$ is calculated as follows,
@@
Researchers from NVIDIA introduced a new SAM variant, NanoSAM [nanosam], based on MobileSAM. It is designed to achieve real-time performance on NVIDIA Jetson Orin platforms using NVIDIA TensorRT. NanoSAM replaces the ViT-based encoder with a pure convolutional network, specifically ResNet18 [he2016deep], while retaining the other components from MobileSAM. NanoSAM is distilled from MobileSAM, and both models are retrained with TensorRT for optimized performance. The image encoder of MobileSAM is optimized with FP32 precision, whereas NanoSAM’s image encoder is built using FP16 precision for faster execution. Inference latency results on Jetson Orin Nano and Jetson AGX Orin demonstrate that NanoSAM is 5x faster than MobileSAM, with minimal accuracy loss. Very recently, a similar work, which builds up an ultra light-weight variant based on the U-Net architecture, named PicoSAM2 [bonazzi2025picosam2], has successfully deployed SAM on Sony IMX500 (a camera sensor with 8MB memory limit.
@@
Due to MobileSAM with TinyViT as backbone still encounters challenges on resource-constrained devices, Wang et al. [wang2023repvit] developed an novel efficient SAM variant, RepViT-SAM, using their newly proposed CNN-based backbone, RepViT [wang2024repvit], as the image encoder. The core idea behind RepViT is to integrate the effective design principles of efficient Vision Transformers (ViTs) into lightweight CNNs. These design principles are applied at three levels: block-level, macro-level, and micro-level. At the block level, RepViT separates the token mixer and channel mixer [yu2023metaformer], reduces the expansion ratio [graham2021levit], and increases the width of blocks. For the macro design, it incorporates early convolutions [xiao2021early] as the input stem, deepens the down-sampling layers [liu2022convnet], employs a simpler classifier, and adjusts the block ratios across stages [radosavovic2020designing]. At the micro level, only 3x3 convolutions are used, and Squeeze-and-Excitation layers [hu2018squeeze] are applied only in odd-numbered blocks. RepViT-SAM is trained by following the same distillation pipeline in [zhang2023faster], claiming to achieve a 10x increase in inference speed compared to MobileSAM.
@@
Figure 10: The framework of distilling EdgeSAM [zhou2023edgesam]. It consists of the encoder-only distillation and the prompt-in-the-loop distillation.
@@
Concurrent with the development of RepViT-SAM, Zhou et al. [zhou2023edgesam] observe that MobileSAM [zhang2023faster] still struggles to achieve real-time performance when deployed on edge devices, such as mobile phones. To address this, they introduced EdgeSAM, which replaces the Transformer-based encoder with the more lightweight and efficient, pure CNN-based RepViT, aiming for better performance on resource-constrained devices. Similar to the approach in [Zhao2023ESAM], Zhou et al. argue that encoder-only distillation is inadequate, as it is task-agnostic and does not fully capture the model’s task-specific needs. To overcome this, they propose the Prompt-in-the-Loop Distillation method, which adds additional supervision to the output masks. The ”Prompt-in-the-Loop” refers to a dynamic sampling strategy that iteratively samples new prompts from the non-overlapping zones of the teacher’s and student’s predicted masks. After several iterations, the accumulated loss is backpropagated to update both the encoder and decoder. To further enhance the output quality, EdgeSAM offers an optional module that embeds granularity priors from specific datasets. The overall framework for distilling EdgeSAM is illustrated in Fig. 10.
@@
To further improve the efficiency of backbone while maintaining strong capability of feature extraction, Zhang et al. [zhang2024efficientvit] introduced EfficientViT [cai2022efficientvit] as image encoder, building up the powerful variant, EfficientViT-SAM. The primary advantage of EfficientViT is its use of the ReLU linear attention mechanism, which facilitates global information interaction while enhancing hardware efficiency. By eliminating the hardware-unfriendly softmax operations and replacing them with ReLU, the attention calculation is reformulated as follows,
@@
Following the similar idea of applying high-efficiency attention mechanisms, Shen et al. [shen2024fastsam3d] proposed FastSAM3D, an efficient segment-anything model specifically designed for 3D volumetric medical images. The key contribution of this work is the development of the 3D Sparse Flash Attention mechanism. This novel attention approach combines the advantages of the 3D Dilated Attention [ding2023longnet], which extends the receptive field, with FlashAttention [dao2022flashattention] to accelerate computations. FastSAM3D utilizes a modified ViT-Tiny as the image encoder, distilled from the ViT-Base encoder, ensuring efficiency without compromising performance. The authors implemented a layer-wise progressive distillation strategy to iteratively align feature maps between the two encoders, as shown in Fig. 11.
Figure 11: The framework of layer-wise progressive distillation of FastSAM3D [shen2024fastsam3d]. It takes the ViT-B as teacher model to distill the efficient image encoder, a 6-layer ViT-Tiny variant.
Building on the approach of FastSAM3D, Song et al. [songa2024sam] introduced a 2D variant called SAM-Lightening. Like its 3D counterpart, SAM-Lightening combines Sparse/Dilated Attention and FlashAttention as a replacement for the standard attention mechanism, while retaining the same image encoder. The key difference lies in the Knowledge Distillation (KD) strategy, referred to as Dynamic Layer-Wise Distillation (DLD). In DLD, a series of time-varying weights $\in[0,1]$ are applied to determine which layers need to be updated and how much each layer contributes to the update process. As training progresses, the entire architecture is gradually optimized. SAM-Lightening reportedly achieves a 30× acceleration compared to SAM-H. The overall distillation framework is similar to FastSAM3D, with the main change being the substitution of the distillation strategy for dynamic layer-wise distillation.
A recent work by Yuan et al. [yuan2024mamba], RWKV-SAM, represents a significant advancement in accelerating SAM by introducing the popular linear attention model [gu2023mamba, peng2023rwkv] as an efficient backbone. In their study, they compare RWKV-based and Mamba-based architectures and select the RWKV-based approach to construct a lightweight version of SAM. The backbone is a hybrid design, with the first two stages consisting of Mobile Convolution Blocks from [sandler2018mobilenetv2], and the final stage built using Vision RWKV blocks [duan2024vision]. More details on RWKV can be found in Section 2.2.2. Additionally, a refinement module is incorporated into the SAM-like architecture to enhance mask quality by fusing features from different levels generated at each stage. The model is trained using a ”distillation-finetune” strategy, where knowledge from SAM-H is first distilled into the backbone, followed by fine-tuning of the entire model. RWKV-SAM demonstrates significant improvements in efficiency, while maintaining comparable segmentation performance to SAM.
@@ 3.1.3 Quantization based Methods
As detailed in previous section, TinySAM [shu2023tinysam] leverages the Hard Mining Full-Stage Distillation to improve the effectiveness of transferring knowledge. To further contract the scale of TinySAM, Shu et al. adopt a post-training quantization to the encoder of TinySAM and the quantified version is named as Q-TinySAM. They take the quantization basically following the instructions in [yuan2022ptq4vit]. For the matrix multiplication in ViT, $O=AB$ would be quantified to $\hat{O}=\hat{A}\hat{B}$, with scaling factors $s_{A},s_{B}$. The researchers takes a alternative and iterative search strategy for the best $s_{A},s_{B}$ which can minimize the distance between $O$ and $\hat{O}$. The hessian guided metric is used to measure the distance which is approximately calculated as follows:
@@
Instead of taking quantization in an encoder-only manner used for Q-TinySAM, Lv et al. [lv2024ptq4sam] propose a novel framework that can directly take post-training quantization on SAM, namely PTQ4SAM. They start the work by first discovering the two challenges after traditional PTQ: 1) the existence of bimodal distribution which has negative effects on quantization quality and 2) the obvious discrepancy in the distribution of different attention mechanisms. Therefore, researchers bring up two strategies to solve them respectively: the Bimodal Integration (BIG) and the Adaptive Granularity Quantization (AGQ). For the Bimodal Integration, a sign factor $\gamma$ is introduced to convert the bimodal distribution into a normal distribution. For the Adaptive Granularity Quantization, the key is to adopt an adaptive parameter $\tau$ to adjust the base in the Log2 quantizer [miyashita2016convolutional]. With the use of the AGQ strategy, the discrepancy between post-softmax distributions of different attention is expected to narrow down effectively. The PTQ4SAM is a plug-and-play SAM variant that can be easily deployed to downstream tasks.
Concurrent with PTQ4SAM, a counterpart named PQ-SAM [liu2024pq], further exploits the post-training quantization method customized for SAM’s activations, achieving the first usable 4-bit PTQ for SAM. The core idea lies in transferring SAM’s highly asymmetric activation distribution with extreme outliers into quantization-friendly distribution, for which the Grouped Activation Distribution Transformation (GAPT) is proposed. GAPT firstly adopts tensor-wise outlier truncating to eliminate outliers in activation values, followed by the channel-wise outlier grouping to classify channels of truncated tensors into clusters based on their range. Pairs of learnable shifting and scaling sizes are subsequently allocated to each channel group, reshaping the distribution as follows:
@@ 3.1.4 Pruning based Methods
Figure 12: The overall pipeline of the alternate slimming strategy for SlimSAM [chen20230]. It takes two steps to contract the heavy image encoder of SAM into a light-weight one: 1) Embedding pruning and bottleneck aligning; 2) Bottleneck pruning and embedding aligning.
Chen et al. [chen20230] were the first to develop an effective pruning strategy for reducing the size and complexity of SAM, resulting in the model known as SlimSAM. As discussed in Section 2.3.3, pruning algorithms aim to remove redundant parameters either structurally or individually. When applied to SAM’s heavy encoder, the initial step involves estimating the importance of weights and activation values to determine which should be pruned. The core idea behind assessing importance is to evaluate the discrepancy in loss generated with and without the given parameter. SlimSAM introduces the Disturbed Taylor Importance method, which uses the first-order Taylor expansion to approximate the importance of parameters and introduces Gaussian Noise $\mathcal{N}$ to prevent gradients from becoming zero. This process is formulated as follows,
@@
Unlike Chen et al. [chen20230] adopting the pruning and post-alignment strategy to lighten SAM in a straightforward manner, Abebe et al. [abebe2025supersam] incorporates pruning methods with the one-shot Neural Architecture Search [white2023neural] (NAS, an effective technique of automating the design of efficient neural networks), building up a weight-sharing supernetwork of SAM, named SuperSAM. The overall architecture of SuperSAM is illustrated as Fig. 13. SuperSAM works as a search space of SAM-like subnetworks with size 30-70% smaller than original SAM-B, while some of them still maintaining considerable performance are considered as efficient variants of SAM. The supernetwork is trained by sampling and training subnetworks with flexible image encoder, in which the structured pruning is applied to remove insignificant Transformer layers, with further weight reordering and slicing of MLP in remaining layers. Once the SuperSAM is built up, techniques such as OpenTuner [ansel2014opentuner] can be deployed to search for the architecture meeting both efficiency and accuracy requirement. The emerge of SuperSAM might further broaden the domain of investigating efficient SAM variants.
Figure 13: The architecture of SuperSAM [abebe2025supersam] consists of a flexible image encoder and SAM-like prompt-guided mask decoder. Lightweight SAM variants can be built up by sampling appropriate image encoder meeting predefined efficiency and accuracy requirements.